# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ Abstract
With the rapid advancement of large language models (LLMs), aligning policy models with human preferences has become increasingly critical. Direct Preference Optimization (DPO) has emerged as a promising approach for alignment, acting as an RL-free alternative to Reinforcement Learning from Human Feedback (RLHF). Despite DPO’s various advancements and inherent limitations, an in-depth review of these aspects is currently lacking in the literature. In this work, we present a comprehensive review of the challenges and opportunities in DPO, covering theoretical analyses, variants, relevant preference datasets, and applications. Specifically, we categorize recent studies on DPO based on key research questions to provide a thorough understanding of DPO’s current landscape. Additionally, we propose several future research directions to offer insights on model alignment for the research community.
@@ 1 Introduction
Through pre-training on extensive, high-quality corpora using the next token prediction objective with a huge amount of computational costs, Large Language Models (LLMs)  (OpenAI, 2022; Touvron et al., 2023a; OpenAI, 2024a; Jiang et al., 2023) assimilate comprehensive world knowledge into their internal parameters, demonstrating impressive language understanding and generation abilities. Further, LLMs have been extended to accommodate multi-modality inputs, including both language and vision, thereby giving rise to Large Vision Language Models (LVLMs)  (OpenAI, 2023a; Liu et al., 2024a; Team et al., 2023; Bai et al., 2023). These foundation models serve as a versatile solution and have achieved exceptional performance across a broad spectrum of both language and visio-language tasks, marking a significant milestone in the advancement toward artificial general intelligence.
@@
As these foundation models grow larger and more powerful, they are still found grappling with following the user’s instruction (explicit objective) and fulfilling the goals of being Helpful, Honest and Harmless (implicit objective), which attribute to the misaligned next token prediction task used in the pre-training stage (Leike et al., 2018; Askell et al., 2021; OpenAI, 2023b). Therefore, a typical post-training stage, known as preference optimization (e.g., Reinforcement Learning from Human Feedback, RLHF), is additionally performed at the response level to align pre-trained language models with the user’s intentions and ensure they remain helpful, honest, and harmless (Ouyang et al., 2022a; Dai et al., 2024; Sun et al., 2023). RLHF first trains an explicit reward model on collected human preferences. Subsequently, RLHF fine-tunes the policy model (i.e., the LLM targeted for fine-tuning) with reinforcement learning (RL) algorithms (e.g., Proximal Policy Optimization(PPO; (Schulman et al., 2017a))) to output responses which can maximize the response reward rated by the reward model but not deviate too far from a reference model constrained by KL-divergence. Nevertheless, RLHF requires meticulous hyper-parameters tuning and extensive computational resource to maintain the RL training stability. Moreover, some research has identified several challenges associated with this explicit reward modeling, such as reward hacking (Casper et al., 2023), reward misspecification (Pan et al., 2022) and out-of-distribution generalization (Tien et al., 2023).
@@
To avoid the aforementioned limitations of RLHF, a range of RL-free preference optimization methods have been proposed. Yuan et al. (2023); Dong et al. (2023); Liu et al. (2024b); Song et al. (2024) propose sampling multiple responses from the policy model and rating them with a well-trained reward model. Then, instead of using an RL algorithm, the policy model is directly fine-tuned under the supervision of the best-of-$K$ response (known as reject sampling) or by applying a ranking loss to the ranked responses. On the other hand, staring from the KL-constrained reward maximization objective in RL, Direct Preference Optimization (DPO; (Rafailov et al., 2023)) derive its learning objective, specifically a simple maximum likelihood objective, on offline preference data, which is directly formulated over the policy model and a reference model, thereby bypassing the explicit reward model training phase and eliminating the need for reinforcement learning optimization. Indeed, the optimization objective of DPO is equivalent to the Bradley Terry model (Bradley and Terry, 1952a) with a implicit reward function that is parameterized by the policy model itself. Compared to RLHF, DPO has been demonstrated stable, performant, and computationally lightweight in various applications (Rafailov et al., 2023; Ethayarajh et al., 2024; Ivison et al., 2024).
@@
More recently, some studies have indicated that despite avoiding computationally expensive reinforcement learning, DPO still encounters some substantial challenges. For instance, implicit reward modeling in DPO might lead to a biased policy that favors out-of-distribution responses (Xu et al., 2024a; Saeidi et al., 2024), offline DPO is empirically inferior to online alignment methods (Ivison et al., 2024), models that undergo alignment might experience an alignment tax (Lin et al., 2024a; Lu et al., 2024a), et, etc. Consequently, various improved versions of DPO have been proposed recently, including KTO (Ethayarajh et al., 2024), IPO (Azar et al., 2023), CPO (Xu et al., 2024b), ORPO (Hong et al., 2024), simPO (Meng et al., 2024), and others (Lu et al., 2024b; Xiao et al., 2024; Zeng et al., 2024). With the rapid progress in DPO, there is an urgent need for a comprehensive review to help researchers identify emerging trends and challenges in this field. We have observed several concurrent studies on LLM alignment that are relevant to our work (Ji et al., 2023; Wang et al., 2023a; Shen et al., 2023). However, the existing review papers primarily focus on the overall alignment of LLMs, including instruction fine-tuning and RLHF. The sections of these studies related to DPO are insufficient to capture the rapid advancements currently unfolding in this area. Furthermore, these reviews tend to focus on alignment within the context of language models, without providing a thorough introduction to the applications and datasets specific to DPO.
@@
  • Effect of Implicit Reward Modeling. DPO circumvents the need to train an explicit reward model by establishing a direct mapping from reward functions to optimal policies. Consequently, studies have examined the generalization capabilities of the implicit reward modeling employed in DPO (Lin et al., 2024b; Li et al., 2024a; Yang et al., 2024a; Jia, 2024).
@@
  • Effect of KL Penalty Coefficient and Reference Model. Both the optimization objective of the RL and DPO involve a Kullback-Leibler (KL) divergence regularization, which constrains the policy model to remain within a specified proximity to the reference model. Therefore, some recent studies have investigated the impact of the KL penalty coefficient and the choice of the reference model (Liu et al., 2024c; Xu et al., 2024a; Feng et al., 2024; Rafailov et al., 2024a).
@@
  • Effect of Different Feedback. DPO uses point-wise reward and pair-wise preference data to provide reward signals. However, obtaining high-quality pair-wise preference data is both costly and time-consuming, posing challenges for scalability. Additionally, instance-level optimization may not fully leverage the potential of preference data. Therefore, some studies employ other forms of feedback (e.g. List-wise, Binary, Step-wise, Token-wise, etc) as the reward signal for optimization (Dong et al., 2023; Yuan et al., 2023; Ethayarajh et al., 2024; Zeng et al., 2024; Chen et al., 2024a; Xu et al., 2024b).
@@
  • Online DPO. Compared to online RLHF, DPO utilizes pre-collected preference data and is considered an offline preference optimization method. Some studies have highlighted the performance gap between online and offline algorithms (Tang et al., 2024; Wang et al., 2024a). To address this gap, recent research has explored iterative and online variants of DPO, as well as strategies for efficiently collecting new preference datasets (Xu et al., 2024c; Guo et al., 2024a; Yuan et al., 2024a; Chen et al., 2024b).
@@
  • Reward Hacking. Reward hacking is a long-standing problem in RL where the policy achieves a high reward but fails to meet the actual objective (Dubois et al., 2024; Singhal et al., 2023). Recent studies have found reward hacking also exists in both RLHF and DPO regardless the explicit or implicit reward model, which exploits potential shortcuts (e.g., response length and style) to develop specific response patterns to hack the reward model (Kabir et al., 2024; Wang et al., 2023b; Park et al., 2024). To overcome this limitation, some methods have been proposed to avoid such weakness been utilized (Park et al., 2024; Yuan et al., 2024b; Meng et al., 2024; Liu et al., 2024d)
@@
  • Alignment Tax. Preference optimization aims to align models with human preferences. However, prior studies have found a phenomenon known as the alignment tax, which refers to that improvements in alignment objective can lead to a decrease in performance compared to a SFT model (Ouyang et al., 2022a). Consequently, some studies have investigated the alignment tax and proposed methods to reduce its effect (Lin et al., 2024a; Lou et al., 2024a; Guo et al., 2024b).
@@
We hope our review will help researchers capture new trends and challenges in this field, explore the potential of DPO in aligning LLMs and Multi-modal LLMs (MLLM) , and contribute to building a more scalable and generalizable DPO. Specifically, we believe that future research should prioritize the development of more advanced DPO variants that: (i) go beyond instance-level feedback to capture more fine-grained and accurate rewards; (ii) exhibit competitive or superior generalization capabilities compared to online RLHF by leveraging data, learning objectives, and rewards; and (iii) facilitate the development of sophisticated applications, such as deep reasoning systems like OpenAI o1 (OpenAI, 2024b), mixed-modal models like Chamelon (Team, 2024).
@@ 2 Preliminary
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5em,font=,, where level=2text width=10em,font=,, where level=3text width=11em,font=,, where level=4text width=7em,font=,, [ X-DPO, ver [ Theory [ Research Questions (§3) [ Why DPO [ InstructGPT (Ouyang et al., 2022a)DPO (Rafailov et al., 2023) , leaf, text width=35em ] ] [ Generalization Ability of
Implicit Reward Model [ Implicit Reward Analysis (Lin et al., 2024b; Li et al., 2024a)
@@
Generalize Reward Model (Jia, 2024; Yang et al., 2024a; Azar et al., 2023) , leaf, text width=35em ] ] [ Different Feedback [ RAFT (Dong et al., 2023)RRHF (Yuan et al., 2023)UltraInteract (Yuan et al., 2024c)
sDPO (Kim et al., 2024)StepDPO (Lai et al., 2024)
KTO (Ethayarajh et al., 2024) TDPO (Zeng et al., 2024)
@@
SPIN (Chen et al., 2024a) CPO (Xu et al., 2024b)IRPO (Pang et al., 2024) , leaf, text width=35em ] ] [ Reward Hacking [ Length Exploitation Analysis (Dubois et al., 2024)
 (Singhal et al., 2023; Kabir et al., 2024; Wang et al., 2023b)
R-DPO (Park et al., 2024)LIFT-DPO (Yuan et al., 2024b)
SimPO (Meng et al., 2024)LD-DPO (Liu et al., 2024d) , leaf, text width=35em ] ] [ KL penalty and
Reference Model [ Analysis (Rafailov et al., 2024a; Feng et al., 2024)
 (Liu et al., 2024c; Xu et al., 2024a)
$\beta$-DPO (Wu et al., 2024a)CPO (Xu et al., 2024b)ORPO (Hong et al., 2024)
@@
SimPO (Meng et al., 2024) , leaf, text width=35em ] ] [ Online DPO [ Analysis (Tang et al., 2024; Wang et al., 2024a)
OAIF (Guo et al., 2024a)OFS-DPO (Qi et al., 2024)
OPTune (Chen et al., 2024b)Self-rewarding (Yuan et al., 2024d)
@@
PCO (Xu et al., 2024c) , leaf, text width=35em ] ] [ Alignment Tax [ Analysis (Ouyang et al., 2022a; Kim and Seo, 2024; Lu et al., 2024a)
SPO (Lou et al., 2024a)PAFT (Pentyala et al., 2024)
AMA (Lin et al., 2024a)Disperse-then-Merge (Fu et al., 2024)
@@
CPO (Guo et al., 2024b)MODPO (Zhou et al., 2024a) , leaf, text width=35em ] ] ] [ Variants [ KTO (Ethayarajh et al., 2024)CPO (Xu et al., 2024b)IPO (Azar et al., 2023)
@@
SimPO (Meng et al., 2024) , leaf, text width=35em ] ] ] [ Datasets [ Human Labeled (§4.1) [ Summarize From Human Feedback  (Stiennon et al., 2022) Tasksource (Sileo, 2024)
Webgpt (Nakano et al., 2022) HelpSteer (Wang et al., 2023c)
OpenAssistant (Köpf et al., 2023) SHP (Ethayarajh et al., 2022)
HelpSteer2 (Wang et al., 2024b) RLHF-V-Dataset (Yu et al., 2024a)
HH-RLHF (Bai et al., 2022a) Chatbot Arena (Zheng et al., 2023a)
@@
Alpaca Farm Human (Dubois et al., 2024) , leaf, text width=35em ] ] [ AI Labeled (§4.2) [ RLAIF-V (Yu et al., 2024b) Step-DPO (Lai et al., 2024)
UltraFeedback (Cui et al., 2023a) Capbara (Argilla, 2024)
Mlabonne (Malabonne, 2024) RewardBench (Lambert et al., 2024)
Magpie (Xu et al., 2024d) Zerolink (Zerolink, 2024a)
Nectar (Zhu et al., 2023) OrcaSlimOrca (Lian et al., 2023)
Alpaca Farm GPT-4 (Dubois et al., 2024) sqrti/SPA-VL (Zhang et al., 2024a)
MMInstruction/VLFeedback (Li et al., 2023a) Py-dpo(Jondurbin, 2024a)
Truthy-dpo (Zerolink, 2024b) Chatml-dpo-pairs (Mukherjee et al., 2023)
@@
Zsql-postgres-dpo (Zerolink, 2024a) tldr-preference-dpo (davanstrien, 2024) , leaf, text width=35em ] ] ] [ Application [ Language Models(§5.1) [ Reasoning [ Step-DPO (Lai et al., 2024) Step-Controlled DPO (Lu et al., 2024b)
@@
IRPO (Pang et al., 2024) DPO-ST (Wang et al., 2024c) MAPO  (She et al., 2024) , leaf, text width=35em ] ] [ Instruction-following [ AUTOIF (Dong et al., 2024) DeMoRecon (Yang et al., 2024b)
@@
Conifer (Sun et al., 2024) FIPO (Lu et al., 2024c) RS-DPO (Khaki et al., 2024) , leaf, text width=35em ] ] [ Hallucination [ FLAME (Lin et al., 2024c) HALU-J (Wang et al., 2024d)
@@
HaluQuestQA (Sachdeva et al., 2024) Factuality FT (Tian et al., 2023) , leaf, text width=35em ] ] [ Code [ Optimise (Gee et al., 2024) Stable Code (Pinnaparaju et al., 2024)
@@
DPA (Nichols et al., 2024) Instruct-Code-Llama (Liu et al., 2024e) , leaf, text width=35em ] ] ] [ Multi-modal
@@
Models (§5.2) [ Multi-modal understanding [ RLHF-V (Yu et al., 2024a) RLAIF-V (Yu et al., 2024b)POVID (Zhou et al., 2024b)
HSA-DPO (Xiao et al., 2024)VLFeedback (Li et al., 2023a)
@@
LLaVA-Hound (Zhang et al., 2024b)Cat (Ye et al., 2024) , leaf, text width=35em ] ] [ Multi-modal generation [ Diffusion-DPO (Wallace et al., 2023)D3PO (Yang et al., 2024c)
Tango (Majumder et al., 2024)Emo-DPO (Gao et al., 2024)
SpeechAlign (Zhang et al., 2024c)SpeechWorthy (Cho et al., 2024)
@@
Text2Motion  (Sheng et al., 2024) , leaf, text width=35em ] ] ] [ Others (§5.3) [ Recommendation [ Softmax-DPO (Chen et al., 2024c)DMPO (Bai et al., 2024) , leaf, text width=35em ] ] [ Science [ Drug delivery (Hu et al., 2024) ConPro (Nguyen et al., 2024)
Protein-DPO (Widatalla et al., 2024) , leaf, text width=35em ] ] ] ] ]
@@
Figure 1: Taxonomy of research in XDPO that consists of theory analysis, variants, datasets and applications
@@
In this section, we will revisit the foundational concepts of RLHF (Ouyang et al., 2022a) (Reinforcement Learning from Human Feedback) and DPO (Rafailov et al., 2023) (Direct Preference Optimization) to highlight the necessity of DPO fine-tuning as a solution to RLHF’s shortcomings.

### v3

@@ Abstract
With the rapid advancement of large language models (LLMs), aligning policy models with human preferences has become increasingly critical. Direct Preference Optimization (DPO) has emerged as a promising approach for alignment, acting as an RL-free alternative to Reinforcement Learning from Human Feedback (RLHF). Despite DPO’s various advancements and inherent limitations, an in-depth review of these aspects is currently lacking in the literature. In this work, we present a comprehensive review of the challenges and opportunities in DPO, covering theoretical analyses, variants, relevant preference datasets, and applications. Specifically, we categorize recent studies on DPO based on key research questions to provide a thorough understanding of DPO’s current landscape. Additionally, we propose several future research directions to offer insights on model alignment for the research community. An updated collection of relevant papers can be found on https://github.com/Mr-Loevan/DPO-Survey.
@@ 1 Introduction
Through pre-training on extensive, high-quality corpora using the next token prediction objective with a huge amount of computational costs, Large Language Models (LLMs)  OpenAI (2022); Touvron et al. (2023a); OpenAI (2024a); Jiang et al. (2023) assimilate comprehensive world knowledge into their internal parameters, demonstrating impressive language understanding and generation abilities. Further, LLMs have been extended to accommodate multi-modality inputs, including both language and vision, thereby giving rise to Large Vision Language Models (LVLMs)  OpenAI (2023a); Liu et al. (2024a); Team et al. (2023); Bai et al. (2023). These foundation models serve as a versatile solution and have achieved exceptional performance across a broad spectrum of both language and visio-language tasks, marking a significant milestone in the advancement toward artificial general intelligence.
@@
As these foundation models grow larger and more powerful, they are still found grappling with following the user’s instruction (explicit objective) and fulfilling the goals of being Helpful, Honest and Harmless (implicit objective), which attribute to the misaligned next token prediction task used in the pre-training stage Leike et al. (2018); Askell et al. (2021); OpenAI (2023b). Therefore, a typical post-training stage, known as preference optimization (e.g., Reinforcement Learning from Human Feedback, RLHF), is additionally performed at the response level to align pre-trained language models with the user’s intentions and ensure they remain helpful, honest, and harmless Ouyang et al. (2022a); Dai et al. (2024); Sun et al. (2023). RLHF first trains an explicit reward model on collected human preferences. Subsequently, RLHF fine-tunes the policy model (i.e., the LLM targeted for fine-tuning) with reinforcement learning (RL) algorithms (e.g., Proximal Policy Optimization(PPO; Schulman et al. (2017a))) to output responses which can maximize the response reward rated by the reward model but not deviate too far from a reference model constrained by KL-divergence. Nevertheless, RLHF requires meticulous hyper-parameters tuning and extensive computational resource to maintain the RL training stability. Moreover, some research has identified several challenges associated with this explicit reward modeling, such as reward hacking Casper et al. (2023), reward misspecification Pan et al. (2022) and out-of-distribution generalization Tien et al. (2023).
@@
To avoid the aforementioned limitations of RLHF, a range of RL-free preference optimization methods have been proposed. Yuan et al. (2023); Dong et al. (2023); Liu et al. (2024b); Song et al. (2024) propose sampling multiple responses from the policy model and rating them with a well-trained reward model. Then, instead of using an RL algorithm, the policy model is directly fine-tuned under the supervision of the best-of-$K$ response (known as reject sampling) or by applying a ranking loss to the ranked responses. On the other hand, staring from the KL-constrained reward maximization objective in RL, Direct Preference Optimization (DPO; Rafailov et al. (2023)) derive its learning objective, specifically a simple maximum likelihood objective, on offline preference data, which is directly formulated over the policy model and a reference model, thereby bypassing the explicit reward model training phase and eliminating the need for reinforcement learning optimization. Indeed, the optimization objective of DPO is equivalent to the Bradley Terry model Bradley and Terry (1952a) with a implicit reward function that is parameterized by the policy model itself. Compared to RLHF, DPO has been demonstrated stable, performant, and computationally lightweight in various applications Rafailov et al. (2023); Ethayarajh et al. (2024); Ivison et al. (2024).
@@
More recently, some studies have indicated that despite avoiding computationally expensive reinforcement learning, DPO still encounters some substantial challenges. For instance, implicit reward modeling in DPO might lead to a biased policy that favors out-of-distribution responses Xu et al. (2024a); Saeidi et al. (2024), offline DPO is empirically inferior to online alignment methods Ivison et al. (2024), models that undergo alignment might experience an alignment tax Lin et al. (2024a); Lu et al. (2024a), et, etc. Consequently, various improved versions of DPO have been proposed recently, including KTO Ethayarajh et al. (2024), IPO Azar et al. (2023), CPO Xu et al. (2024b), ORPO Hong et al. (2024), simPO Meng et al. (2024), and others Lu et al. (2024b); Xiao et al. (2024); Zeng et al. (2024). With the rapid progress in DPO, there is an urgent need for a comprehensive review to help researchers identify emerging trends and challenges in this field. We have observed several concurrent studies on LLM alignment that are relevant to our work Ji et al. (2023); Wang et al. (2023a); Shen et al. (2023). However, the existing review papers primarily focus on the overall alignment of LLMs, including instruction fine-tuning and RLHF. The sections of these studies related to DPO are insufficient to capture the rapid advancements currently unfolding in this area. Furthermore, these reviews tend to focus on alignment within the context of language models, without providing a thorough introduction to the applications and datasets specific to DPO.
@@
  • Effect of Implicit Reward Modeling. DPO circumvents the need to train an explicit reward model by establishing a direct mapping from reward functions to optimal policies. Consequently, studies have examined the generalization capabilities of the implicit reward modeling employed in DPO Lin et al. (2024b); Li et al. (2024a); Yang et al. (2024a); Jia (2024).
@@
  • Effect of KL Penalty Coefficient and Reference Model. Both the optimization objective of the RL and DPO involve a Kullback-Leibler (KL) divergence regularization, which constrains the policy model to remain within a specified proximity to the reference model. Therefore, some recent studies have investigated the impact of the KL penalty coefficient and the choice of the reference model Liu et al. (2024c); Xu et al. (2024a); Feng et al. (2024); Rafailov et al. (2024a).
@@
  • Effect of Different Feedback. DPO uses point-wise reward and pair-wise preference data to provide reward signals. However, obtaining high-quality pair-wise preference data is both costly and time-consuming, posing challenges for scalability. Additionally, instance-level optimization may not fully leverage the potential of preference data. Therefore, some studies employ other forms of feedback (e.g. List-wise, Binary, Step-wise, Token-wise, etc) as the reward signal for optimization Dong et al. (2023); Yuan et al. (2023); Ethayarajh et al. (2024); Zeng et al. (2024); Chen et al. (2024a); Xu et al. (2024b).
@@
  • Online DPO. Compared to online RLHF, DPO utilizes pre-collected preference data and is considered an offline preference optimization method. Some studies have highlighted the performance gap between online and offline algorithms Tang et al. (2024); Wang et al. (2024a). To address this gap, recent research has explored iterative and online variants of DPO, as well as strategies for efficiently collecting new preference datasets Xu et al. (2024c); Guo et al. (2024a); Yuan et al. (2024a); Chen et al. (2024b).
@@
  • Reward Hacking. Reward hacking is a long-standing problem in RL where the policy achieves a high reward but fails to meet the actual objective Dubois et al. (2024); Singhal et al. (2023). Recent studies have found reward hacking also exists in both RLHF and DPO regardless the explicit or implicit reward model, which exploits potential shortcuts (e.g., response length and style) to develop specific response patterns to hack the reward model Kabir et al. (2024); Wang et al. (2023b); Park et al. (2024). To overcome this limitation, some methods have been proposed to avoid such weakness been utilized Park et al. (2024); Yuan et al. (2024b); Meng et al. (2024); Liu et al. (2024d)
@@
  • Alignment Tax. Preference optimization aims to align models with human preferences. However, prior studies have found a phenomenon known as the alignment tax, which refers to that improvements in alignment objective can lead to a decrease in performance compared to a SFT model Ouyang et al. (2022a). Consequently, some studies have investigated the alignment tax and proposed methods to reduce its effect Lin et al. (2024a); Lou et al. (2024a); Guo et al. (2024b).
@@
We hope our review will help researchers capture new trends and challenges in this field, explore the potential of DPO in aligning LLMs and Multi-modal LLMs (MLLM) , and contribute to building a more scalable and generalizable DPO. Specifically, we believe that future research should prioritize the development of more advanced DPO variants that: (i) go beyond instance-level feedback to capture more fine-grained and accurate rewards; (ii) exhibit competitive or superior generalization capabilities compared to online RLHF by leveraging data, learning objectives, and rewards; and (iii) facilitate the development of sophisticated applications, such as deep reasoning systems like OpenAI o1 OpenAI (2024b), mixed-modal models like Chamelon Team (2024).
@@ 2 Preliminary
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5em,font=,, where level=2text width=10em,font=,, where level=3text width=11em,font=,, where level=4text width=7em,font=,, [ DPO, ver [ Theory [ Research Questions (§3) [ Why DPO [ InstructGPT Ouyang et al. (2022a)DPO Rafailov et al. (2023) , leaf, text width=35em ] ] [ Implicit Reward Modeling [ Implicit Reward Analysis Lin et al. (2024b); Li et al. (2024a); Cen et al. (2025)
@@
Generalize Reward Model Jia (2024); Yang et al. (2024a); Azar et al. (2023) , leaf, text width=35em ] ] [ Different Feedback [ RAFT Dong et al. (2023)RRHF Yuan et al. (2023)UltraInteract Yuan et al. (2024c)
sDPO Kim et al. (2024)StepDPO Lai et al. (2024)
KTO Ethayarajh et al. (2024) TDPO Zeng et al. (2024)M-DPO Xiong et al. (2025)
@@
SPIN Chen et al. (2024a) CPO Xu et al. (2024b)IRPO Pang et al. (2024) , leaf, text width=35em ] ] [ KL penalty and
Reference Model [ Analysis Rafailov et al. (2024a); Feng et al. (2024)
 Liu et al. (2024c); Xu et al. (2024a)
$\beta$-DPO Wu et al. (2024a)CPO Xu et al. (2024b)ORPO Hong et al. (2024)
@@
SimPO Meng et al. (2024)$\epsilon$-DPO Lee et al. (2025)TR-DPO Gorbatovski et al. (2025) , leaf, text width=35em ] ] [ Online DPO [ Analysis Tang et al. (2024); Wang et al. (2024a)
OAIF Guo et al. (2024a)OFS-DPO Qi et al. (2024)
OPTune Chen et al. (2024b)Self-rewarding Yuan et al. (2024d)
PCO Xu et al. (2024c)HPO Bose et al. (2025) , leaf, text width=35em ] ] [ Reward Hacking [ Length Exploitation Analysis Dubois et al. (2024)
 Singhal et al. (2023); Kabir et al. (2024); Wang et al. (2023b)
R-DPO Park et al. (2024)LIFT-DPO Yuan et al. (2024b)SimPO Meng et al. (2024)
@@
LD-DPO Liu et al. (2024d)POWER Rashidinejad and Tian (2025) , leaf, text width=35em ] ] [ Alignment Tax [ Analysis Ouyang et al. (2022a); Kim and Seo (2024); Lu et al. (2024a)
SPO Lou et al. (2024a)PAFT Pentyala et al. (2024)
AMA Lin et al. (2024a)Disperse-then-Merge Fu et al. (2024)
@@
CPO Guo et al. (2024b)MODPO Zhou et al. (2024a)RCFT Xiao et al. (2025) , leaf, text width=35em ] ] ] [ Variants [ KTO Ethayarajh et al. (2024)CPO Xu et al. (2024b)IPO Azar et al. (2023)
@@
SimPO Meng et al. (2024)ORPO Hong et al. (2024) , leaf, text width=35em ] ] ] [ Datasets [ Human Labeled (§4.1) [ Summarize From Human Feedback  Stiennon et al. (2022) Tasksource Sileo (2024)
Webgpt Nakano et al. (2022) HelpSteer Wang et al. (2023c)
OpenAssistant Köpf et al. (2023) SHP Ethayarajh et al. (2022)
HelpSteer2 Wang et al. (2024b) RLHF-V-Dataset Yu et al. (2024a)
HH-RLHF Bai et al. (2022a) Chatbot Arena Zheng et al. (2023a)
Alpaca Farm Human Dubois et al. (2024) HelpSteer3 Wang et al. (2025a)
@@
MM-RLHF Zhang et al. (2025) MMSafe-PO Li et al. (2025a) , leaf, text width=35em ] ] [ AI Labeled (§4.2) [ RLAIF-V Yu et al. (2024b) Step-DPO Lai et al. (2024)
UltraFeedback Cui et al. (2023a) Capbara Argilla (2024)
Mlabonne Malabonne (2024) RewardBench Lambert et al. (2024)
Magpie Xu et al. (2024d) Zerolink Zerolink (2024a)
Nectar Zhu et al. (2023) OrcaSlimOrca Lian et al. (2023)
Alpaca Farm GPT-4 Dubois et al. (2024) sqrti/SPA-VL Zhang et al. (2024a)
MMInstruction/VLFeedback Li et al. (2023a) Py-dpoJondurbin (2024a)
Truthy-dpo Zerolink (2024b) Chatml-dpo-pairs Mukherjee et al. (2023)
Zsql-postgres-dpo Zerolink (2024a) tldr-preference-dpo davanstrien (2024)
RewardBench2 Malik et al. (2025) VL-RewardBench Li et al. (2025b)
CodeSteer-DPO-Dataset Chen et al. (2025) ANAH Gu et al. (2025)
@@
Unified Preference Dataset Wang et al. (2025b) MM-IFDPO-23K Ding et al. (2025) , leaf, text width=35em ] ] ] [ Application [ Language Models(§5.1) [ Reasoning [ Step-DPO Lai et al. (2024) Step-Controlled DPO Lu et al. (2024b)
@@
IRPO Pang et al. (2024) DPO-ST Wang et al. (2024c) MAPO  She et al. (2024) , leaf, text width=35em ] ] [ Instruction-following [ AUTOIF Dong et al. (2024) DeMoRecon Yang et al. (2024b)
@@
Conifer Sun et al. (2024) FIPO Lu et al. (2024c) RS-DPO Khaki et al. (2024) , leaf, text width=35em ] ] [ Hallucination [ FLAME Lin et al. (2024c) HALU-J Wang et al. (2024d)
@@
HaluQuestQA Sachdeva et al. (2024) Factuality FT Tian et al. (2023) , leaf, text width=35em ] ] [ Code [ Optimise Gee et al. (2024) Stable Code Pinnaparaju et al. (2024)
@@
DPA Nichols et al. (2024) Instruct-Code-Llama Liu et al. (2024e) , leaf, text width=35em ] ] ] [ Multi-modal
@@
Models (§5.2) [ Multi-modal understanding [ RLHF-V Yu et al. (2024a) RLAIF-V Yu et al. (2024b)POVID Zhou et al. (2024b)
HSA-DPO Xiao et al. (2024)VLFeedback Li et al. (2023a)
@@
LLaVA-Hound Zhang et al. (2024b)Cat Ye et al. (2024) , leaf, text width=35em ] ] [ Multi-modal generation [ Diffusion-DPO Wallace et al. (2023)D3PO Yang et al. (2024c)
Tango Majumder et al. (2024)Emo-DPO Gao et al. (2024)
SpeechAlign Zhang et al. (2024c)SpeechWorthy Cho et al. (2024)
@@
Text2Motion  Sheng et al. (2024) , leaf, text width=35em ] ] ] [ Others (§5.3) [ Recommendation [ Softmax-DPO Chen et al. (2024c)DMPO Bai et al. (2024) , leaf, text width=35em ] ] [ Science [ Drug delivery Hu et al. (2024) ConPro Nguyen et al. (2024)
Protein-DPO Widatalla et al. (2024) , leaf, text width=35em ] ] ] ] ]
@@
Figure 1: Taxonomy of research in DPO that consists of theory analysis, variants, datasets and applications
@@
In this section, we will revisit the foundational concepts of RLHF Ouyang et al. (2022a) (Reinforcement Learning from Human Feedback) and DPO Rafailov et al. (2023) (Direct Preference Optimization) to highlight the necessity of DPO fine-tuning as a solution to RLHF’s shortcomings.